OpenStudy (anonymous):

What are the zero's of this quadratic? 3x^2=4800

OpenStudy (anonymous):

\[3x ^{2}=4800\]

jimthompson5910 (jim_thompson5910):

Hint: 3x^2 = 4800 x^2 = 4800/3 ... Divide both sides by 3. x^2 = 1600

OpenStudy (anonymous):

is it x= +-40

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thanks I was just checking my work (: